OpenStudy (anonymous):

Calculus and Vectors ---> If a point P lies on y-axis, what are the direction angles of OP ? & The vector u has a magnitude of 10. The x- and y-components of u are positive and equal and its z-component is 4. Determine x and ycomponent of vector u . Step by step please!

OpenStudy (anonymous):

its 90,0,90 (all in degrees)

OpenStudy (anonymous):

let the vector be <a,a,4> its agnitude is given by sqrt(a^2 +a^2 +4^2)=10 or 2a^2+16 =100 or 2a^2= 84 or a^2 84/2=42 hence a= +sqrt(42)